OpenStudy (anonymous):

Make a table of values for y=-|2x+5|

OpenStudy (zpupster):

y=f(x) so we will use 2,1,0,-1,-2 for x i will do one as an example: f(2) = -|2(2) +5| = -(4+5) = -(9) = -9 in your table x value =2 y value =-9 f(1)= -|2(1) +5| f(-0)=-|2(0) +5| f(-1)=-|2(-1) +5| f(-2)=-|2(-2) +5|